OpenStudy (anonymous):

If a 40.0 kg object reduces its speed from 30 m/s to 15 m/s in 10 seconds, . What is the formula that could be used ?

OpenStudy (anonymous):

for the acceleration?

OpenStudy (mr.math):

Or for the force?

OpenStudy (mr.math):

You have acceleration a=(15-30)/10=-1.5 m/s^2.

OpenStudy (mr.math):

The force F=ma=40*(-1.5)=-60N

OpenStudy (anonymous):

The force. Can you show the steps too please ? Thanks .

OpenStudy (mr.math):

There's only one step, which is substitution into the formula F=ma, where m is mass and a is acceleration.

OpenStudy (anonymous):

Thanks . I just needed that part .

OpenStudy (mr.math):

You're welcome.